OpenStudy (superhelp101):

Choose the point that lies on the graph of the piecewise function. f of x equals negative the quantity of x plus 5 end quantity squared if x is less than or equal to negative 3 and equals 2x minus 5 if x is greater than negative (-2, 9) (-3, -11) (-7, -4) (-4, -13)

OpenStudy (superhelp101):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

can you draw out the piecewise function?

OpenStudy (superhelp101):

yes I will try..

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

(-2, 9) means x = -2, y = 9

jimthompson5910 (jim_thompson5910):

when x = -2, what inequality is being satisfied?

OpenStudy (superhelp101):

x greater than -3

jimthompson5910 (jim_thompson5910):

x > -3 -2 > -3 ... replace x with -2 -2 > -3 is true, so you are correct

jimthompson5910 (jim_thompson5910):

so you pick on the piece that corresponds to x > -3

jimthompson5910 (jim_thompson5910):

that's the second piece, therefore y = 2x-5 when x > -3

jimthompson5910 (jim_thompson5910):

and you plug in x = -2 y = 2x - 5 y = 2(-2) - 5 y = -4 - 5 y = -9 we want +9 and NOT -9 So the point (-2,9) is NOT on the curve of this piecewise function

jimthompson5910 (jim_thompson5910):

see how this process works?

OpenStudy (superhelp101):

yes I will try to find it and i will get back in a sec.. thx :)

OpenStudy (superhelp101):

It's B right?

OpenStudy (superhelp101):

wait or D

OpenStudy (superhelp101):

cuz I plugged in y=2(-3)-5 y=-11

jimthompson5910 (jim_thompson5910):

wait, what is "y=2(-3)-5 y=-11" in reference to?

OpenStudy (superhelp101):

option 2

jimthompson5910 (jim_thompson5910):

here is a graph of the piecewise function

jimthompson5910 (jim_thompson5910):

pay very close attention to the endpoints (the open and closed circles)

jimthompson5910 (jim_thompson5910):

now let's plot the points on this piecewise function

jimthompson5910 (jim_thompson5910):

points A and D are clearly not on the graph, so they are out B looks like it is on the graph, but it's actually in a hole, which is NOT part of the graph only point C is on the graph

OpenStudy (superhelp101):

oh okay. how did you graph this what did you use

jimthompson5910 (jim_thompson5910):

when x = -7, x <=-3 is true y = -(x+5)^2 -4 = -(-7+5)^2 ... plug in (x,y) = (-7,-4) -4 = -(-2)^2 -4 = -4 ... answer checks out

jimthompson5910 (jim_thompson5910):

geogebra http://www.geogebra.org/cms/en/

OpenStudy (superhelp101):

Do i need to download it?

jimthompson5910 (jim_thompson5910):

I think you can download it or you can use the web-based version. So it's up to you which option works best. I prefer to download and install it so you can use it offline

OpenStudy (superhelp101):

oh okay, thanks! how do I do the wed-based version may I ask for a link for that

jimthompson5910 (jim_thompson5910):

try this out http://web.geogebra.org/chromeapp/

OpenStudy (superhelp101):

oh okay thank you so much! You've been most helpful! :))

jimthompson5910 (jim_thompson5910):

you're welcome